-5d-24=-4(d+6)-d
-5d-24=-4d+-24-d
-4d-24=-4d-24-d
1d+24=+24-d
1d=1d NOT ON MY ANSWER SHEET
a. -5
b. no solution
c.all numbers
d. -24 Thank you

Answer by Edwin McCravy(20054)

Maybe do it this slightly different way:

-5d - 24 = -4(d + 6) - d

-5d - 24 = -4d - 24 - d

-5d - 24 = -5d - 24 

-5d + 5d = -24 + 24

      0d = 0

You can substitute any number for d in
that equation and it will always be a 
solution since zero times any number
always equals 0.

So the anwer is "c. all numbers".

--------------------

But you could do it your way.

From your result 

  1d = 1d

you could tell that no matter what you
substituted for d, it would always check.

For instance if you substitute (2) for d

  1d = 1d
1(2) = 1(2)
   2 = 2

That checks.

If you substitute (-3) for d

  1d = 1d
1(-3) = 1(-3)
  -3 = -3

That checks.

You have the same thing on both sides of the equation
with  your

1d = 1d 

So any number you substitute on either side will have
to give you the same result on the other side.

Edwin